INFRARED SPECTRA OF ACETYLENE-NITROUS OXIDE TRIMERS: THE (N_{2}O)_{2}-C_{2}H_{2} AND (N_{2}O)_{2}- C_{2}D_{2}

M. DEHGHANY, MAHIN AFSHARI, J. N. OLIAEE, \underline{N. MOAZZEN-AHMADI}, Department of Physics and Astronomy, University of Calgary, Calgary, AB T2N 1N4, CANADA; A.R.W. MCKEL-LAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, CANADA.

The spectra of the (N_{2}O)_{2}-C_{2}H_{2} and ((N_{2}O)_{2}-C_{2}D_{2} trimers in the region of the N_{2}O ν_{1} fundamental ( 2224 cm-1) are observed using a tuneable diode laser to probe a pulsed supersonic slit jet expansion. The observed bands have c- and b-type rotational structure and the trimer has C_{2} symmetry. The structure of the trimer can be thought of as a nonpolar N_{2}O dimer with a C_{2}H_{2} monomer lying above the dimer plane. Search for the other possible isomers of the trimer is currently underway.